### v1

# Identifiability of the Optimal Transport Cost on Finite Spaces

Abstract
The goal of optimal transport (OT) is to find optimal assignments or matchings between data sets which minimize the total cost for a given cost function. However, sometimes the cost function is unknown but we have access to (parts of) the solution to the OT problem, e.g. the OT plan or the value of the objective function. Recovering the cost from such information is called inverse OT and has become recently of certain interest triggered by novel applications, e.g. in social science and economics. This raises the issue under which circumstances such cost is identifiable, i.e., it can be uniquely recovered from other OT quantities. In this work we provide sufficient and necessary conditions for the identifiability of the cost function on finite ground spaces. We find that such conditions correspond to the combinatorial structure of the corresponding linear program.
Keywords: inverse optimal transport, identifiability, cost function, linear programming
MSC 2020 subject classification: Primary 90C08; 52B12; 15A29; secondary 62J07; 62F12
1 Introduction
Numerous natural or social phenomena are often described through the minimization (or maximization) of a cost (or utility) function. In economics, for instance, consumers tend to select products that maximize their utility function (see [GalichonBook]). The utility of a product might depend on ease of access, e.g. distance to the shop it is sold at, and its price and quantity there. Besides, minimization of the total cost, e.g. encoding the distance on a graph or network, can provide understanding of the organization of biological or physical systems, see e.g. [Schiebinger2019, Bunne2023, Bunne2024, Naas2024, Cang2023, Qu2024]. In fact, recently, the applications of optimal transport theory have significantly expanded, driven by the development of faster computational algorithms that enable the handling of larger datasets. This advancement, in turn, has facilitated its broader use in data science including statistics and machine learning, see e.g. [Peyre2018, Kuhn2019, Montesuma2024, Panaretos2020, Sommerfeld2018] and the references given there.
Often these problems can be conveniently mathematically formulated in the optimal transport (OT) framework: Move mass from the target distribution $\mu$ to the source distribution $\nu$ (which are assumed here to have the same total mass) with the least amount of total effort according to some cost function $c$. To be more precise, for $N\in\mathbb{N}$ denote with $\Delta_{N}\coloneqq\{x\in\mathbb{R}^{N}_{+}:\sum_{i=1}^{N}x_{i}=1\}$ the $N$-dimensional probability simplex, where, $\mathbb{R}_{+}$ stands for the set of nonnegative real numbers. Given a cost matrix $c\in\mathbb{R}^{N\times M}$, the OT problem between two probability vectors $\mu\in\Delta_{N}$, $\nu\in\Delta_{M}$ is then defined as
$\mathcal{T}_{c}(\mu,\nu)\coloneqq\min_{\pi\in\Pi(\mu,\nu)}\langle{c},{\pi}\rangle\,,$ (1)
where
[…]
is the Frobenius inner product and $\Pi(\mu,\nu)$ represents the set of probability matrices of size $N\times M$ with marginals $\mu$ and $\nu$, i.e.,
$\Pi(\mu,\nu)\coloneqq\biggl{\{}\pi\in\mathbb{R}^{N\times M}_{+}:\sum_{s=1}^{M}\pi_{i,s}=\mu_{i},\,\sum_{r=1}^{N}\pi_{r,j}=\nu_{j}\text{ for all }i\in\llbracket N\rrbracket,\,j\in\llbracket M\rrbracket\biggr{\}},$
for $\llbracket N\rrbracket\coloneqq\{1,\ldots,N\}$. We call any minimizer $\pi$ of (1) an OT plan between $\mu$ and $\nu$. The value $\mathcal{T}_{c}(\mu,\nu)$ in (1) gives the total cost of the OT problem. Furthermore, we have the dual formulation (cf. [Luenberger2008])
$\mathcal{T}_{c}(\mu,\nu)=\max_{(f,g)\in\Phi_{c}}\langle{f},{\mu}\rangle+\langle{g},{\nu}\rangle\,,$ (2)
where $\Phi_{c}\coloneqq\{(f,g)\in\mathbb{R}^{N}\times\mathbb{R}^{M}\mid f_{i}+g_{j}\leq c_{i,j}\text{ for all }i\in\llbracket N\rrbracket,\,j\in\llbracket M\rrbracket\}$. The elements of a maximizing pair of (2) are called optimal potentials.
In specific settings, the OT plan can represent an optimal assignment w.r.t. the cost $c$. In the discussed economical models, it could represent the optimal assignment of workers to firms or buyers to sellers. The theory of optimal transport on finite ground spaces is classic since the pioneering work of Kantorovitch [Kantorovitch1942, Kantorovitch1958] and can be cast in the framework of linear programming [Bertsimas1997, Luenberger2008]. See also [Brezis2018] for an elementary proof of (2), and the references given there. Once the cost function is specified, the OT plan and total cost can be computed (numerically) from the marginals $\mu$ and $\nu$, see [Peyre2018] for various methods.
However, in many instances, the cost function is (partially) unknown and must be inferred from the observed optimal matching. This is known as inverse optimal transport and was introduced by [InverseSIam] to model migration flows. Since then, inverse OT received great attention, see e.g. [Dupuy2016EstimatingMA, InverseJMLR2023, InversePMLR, InverseSIam, InverseCarlier, andrade2023sparsistency], where further methods for estimation of the cost matrix are provided. However, to put such methods on solid mathematical grounds, it is key to understand when the observed data uniquely determines the cost matrix, which is the goal of this work. That is, we describe conditions on the data so that the cost matrix is identifiable, a terminology we borrowed from the theory of statistical models, where it has a long history (see e.g. [Teicher1963, Lindsay1995, Koopmans1950]). We focus on finitely supported probability measures, which complements our previous work [gonzalezsanz2024nonlinearinverseoptimaltransport] in the continuous case.
Our observations can be described as follows. We observe $K\in\mathbb{N}$ pairs of marginals $(\mu^{(k)},\nu^{(k)})\in\Delta_{N}\times\Delta_{M}$, $k\in\llbracket K\rrbracket$, and (a subset of) the OT information w.r.t. the underlying cost matrix $c$:
  1. (i)
    The corresponding total costs $\alpha^{(k)}\in\mathbb{R}$, $k\in\llbracket K\rrbracket$,
  2. (ii)
    OT plans $\pi^{(k)}\in\Pi(\mu^{(k)},\nu^{(k)})$, $k\in\llbracket K\rrbracket$, and
  3. (iii)
    optimal potentials $(f^{(k)},g^{(k)})\in\Phi_{c}$, $k\in\llbracket K\rrbracket$, such that
$\alpha^{(k)}=\mathcal{T}_{c}(\mu^{(k)},\nu^{(k)})=\langle{c},{\pi^{(k)}}\rangle=\langle{f^{(k)}},{\mu^{(k)}}\rangle+\langle{g^{(k)}},{\nu^{(k)}}\rangle\,.$
Our aim is to recover the cost matrix $c$ from (a combination of) the quantities in i-iii, which are either observed directly or are corrupted by noise. We call $c$ identifiable if it is uniquely determined by (OBS), i.e., if (a subset of) the system $\{\mu^{(k)},\nu^{(k)},\alpha^{(k)},\pi^{(k)},f^{(k)},g^{(k)}\}_{k=1}^{K}$ uniquely determines $c$. Furthermore, we say that $c$ is consistent with the observed OT information if (OBS) holds. Note that, when considering identifiability, we assume that the underlying cost matrix $c$ exists, i.e., there is at least one cost matrix that is consistent with the observed OT information.
It is instructive to distinguish our notion of “inverse optimal transport” from inverse linear programming. Recall that the forward OT problem, i.e., the computation of the total cost $\mathcal{T}_{c}(\mu,\nu)$ and a corresponding OT plan $\pi$ in (1) on finite spaces, amounts to solving a linear program (LP), i.e., for $A\in\mathbb{R}^{E\times D}$, $\bm{c}\in\mathbb{R}^{D}$ and $\bm{b}\in\mathbb{R}^{E}$ the minimization problem
$\min_{\bm{x}\in P(\bm{b})}\langle{\bm{c}},{\bm{x}}\rangle\qquad\text{ where }\quad P(\bm{b})\coloneqq\{\bm{x}\in\mathbb{R}^{D}:A\bm{x}=\bm{b}\,,\bm{x}\geq 0\}\,.$
In contrast, given an observation $\bm{x}^{\prime}\in\mathbb{R}^{D}$ and prior information $\bm{c}^{\prime}\in\mathbb{R}^{D}$, the goal of inverse LP is to solve the following minimization problem
$\min_{\bm{c}}\{\lVert\bm{c}-\bm{c}^{\prime}\rVert:\bm{x}^{\prime}\in\operatorname*{arg\,min}_{\bm{z}\in P(\bm{b})}\langle{\bm{c}},{\bm{z}}\rangle\}\,,$ (3)
where
delimited-∥∥
[…]
is some (semi-)norm on $\mathbb{R}^{D}$. Hence, inverse LP tries to find a $\bm{c}$ close to $\bm{c}^{\prime}$ such that $\bm{x}^{\prime}$ is an optimal solution to the corresponding LP whereas inverse OT aims to recover $c$ from (a combination of) quantities in i-iii. In particular, when the LP is the OT problem, then (3) corresponds to the inverse OT problem in the case that only one OT plan ii is observed, and the objective is made constant (instead of $c\mapsto\lVert c-c^{\prime}\rVert$). The inverse LP problem has been intensively studied in the optimization literature, see e.g. [DempeInverLP, Chan2023, Ahuja2001, Iyengar2005, JianzhongZhang1996, Zhang1999] and references therein. As for inverse OT, most of the work focuses on algorithmic methods to obtain a solution $\bm{c}$ to (3). To put inverse LP into the perspective of the present work, note that identifiability of a solution to the inverse LP problem (3) does not hold in general. Furthermore, inverse LP assumes the observation of a single solution of the LP. Our paradigm is different and is motivated by the notion of inverse problems, where a forward operator (matrix) has to be inverted.
Note that we present our results in the OT framework, but they can also be formulated for more general linear programs. We discuss this in 2.4.
1.1 Organization of the Work
The rest of the paper is organized as follows: In Section 2, we present our main results. This section is divided into subsections, each of which analyzes the identifiability of the cost function under the assumption of observing different OT related quantities in i-iii. They are organized such that gradually information is added such that identifiability becomes easier. In Subsection 2.1, we introduce the notation used throughout the paper, as well as some general results that will be utilized later. The main result of Subsection 2.2 is Theorem 2.6, which provides necessary and sufficient conditions for identifiability when only the total costs i are observed. In Subsection 2.3, we derive the necessary and sufficient conditions for identifiability when only the potentials iii are observed (see Theorem 2.13). Then, Theorem 2.17 in Subsection 2.4 states the identifiability conditions in the setting of only observing the transport plans ii. Subsection 2.5 states necessary and sufficient conditions under the observation of both total costs i and transport plans ii (Theorem 2.21). Finally, the case of full information i-iii is covered by Theorem 2.29 in Subsection 2.6, where necessary and sufficient conditions are given.
In some cases, we may have prior information about the cost function’s structure. An important example is the symmetry of the cost matrix. This will be addressed in Subsection 2.7, where we give identifiability conditions under the observation of total costs and OT plans (see Theorem 2.31) and under the exclusive observation of OT plans (see Theorem 2.32).
As an application of our theory, in Section 3 we propose an estimator for the cost function and derive consistency results in a statistical setting, as well as asymptotic confidence intervals, see Theorem 3.1. Theorem 3.3 is specifically tailored to sparse costs. To this end we introduce a modified notion of sparsity that is adapted to OT and that enjoys corresponding results to statistical sparse recovery.
2 Main Results. Identifiability of the Cost Matrix
This section contains the main results of this paper. In each subsection we derive necessary and sufficient conditions for the identifiability of the cost matrix $c$ considering the different possible combinations of the available information. We refer to Subsection 1.1 for the organization of the different subsections. Note that in all the cases the marginals are assumed to be known (e.g. implicitly given by the knowledge of OT plans).
2.1 Notation and Preliminary Results
As the OT problem (1) is an instance of a linear program (LP), we can make use of the geometry of LPs to achieve identifiability of the cost matrix $c$. To this end, we first introduce the needed tools and terminology from linear programming [Bertsimas1997, Luenberger2008].
Let $(\mu,\nu)\in\Delta_{N}\times\Delta_{M}$ be a pair of marginals and $c\in\mathbb{R}^{N\times M}$ a cost matrix. We denote the by column-major order vectorized version of $c\in\mathbb{R}^{N\times M}$ by
$\bm{c}\coloneqq(\bm{c}_{1},\dots,\bm{c}_{NM})^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\coloneqq(c_{1,1},\dots,c_{N,1},\dots,c_{1,M},\dots,c_{N,M})^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\in\mathbb{R}^{NM}\,.$
The set of transport plans $\Pi(\mu,\nu)$ is a (bounded) polytope. Furthermore, denote the non-empty and compact set of OT plans with
$\operatorname{OTP}_{c}(\mu,\nu)\coloneqq\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}\langle{c},{\pi}\rangle\,.$
A subset $F\subseteq\Pi(\mu,\nu)$ is called a face if there exists a cost matrix $c$ such that $F=\operatorname{OTP}_{c}(\mu,\nu)$. An element $\pi\in\Pi(\mu,\nu)$ is called an extreme point (or vertex) if $\{\pi\}$ is a face or equivalently, if $\pi$ is not a convex combination of two different points in $\Pi(\mu,\nu)$. Denote the set of extreme points of $\Pi(\mu,\nu)$ with $\operatorname{EP}(\mu,\nu)$. As a shortcut, we may write for $\pi\in\Pi(\mu,\nu)$ that $\Pi(\pi)\coloneqq\Pi(\mu,\nu)$ and $\operatorname{EP}(\pi)\coloneqq\operatorname{EP}(\mu,\nu)$ as well as $\operatorname{OTP}_{c}(\pi)=\operatorname{OTP}_{c}(\mu,\nu)$. It follows that every face $F$ of $\Pi(\mu,\nu)$, including $\Pi(\mu,\nu)$ itself, is the convex hull of a unique set of extreme points, i.e., there exist $u^{(1)},\ldots,u^{(L_{F})}\in\operatorname{EP}(\mu,\nu)$ such that
$F=\operatorname{coh}\{u^{(\ell)}:\ell\in\llbracket L_{F}\rrbracket\}\coloneqq\left\{\sum_{\ell=1}^{L_{F}}\lambda_{\ell}\,u^{(\ell)}:\lambda\in\Delta_{L_{F}}\right\}\,.$
Here, $(L_{F}-1)$ is the dimension of the face $F$. In particular, when observing the OT plans ii we can find the faces with minimal dimension that contain them. These faces must then be contained in the set of optimizers. We summarize this observation for later use in the following:
Proposition (RI).
Suppose that we observe $K$ pairs of marginals and the corresponding OT plans ii. Then, for each $k\in\llbracket K\rrbracket$, there exist $L_{k}\in\mathbb{N}$ and $u^{(1,k)},\ldots,u^{(L_{k},k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ such that
$\pi^{(k)}\in\left\{\sum_{\ell=1}^{L_{k}}\lambda_{\ell}u^{(\ell,k)}:\lambda\in\Delta_{L_{k}},\lambda>0\right\}\quad\text{ and }\quad\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}\subseteq\operatorname{OTP}_{c}(\pi^{(k)})\,.$
Remark 2.1 (Information content of (RI)).
(RI) shows that $\pi^{(k)}$ carries more information about $\operatorname{OTP}_{c}(\mu^{(k)},\nu^{(k)})$ for bigger $L_{k}$. The least informative case is $L_{k}=1$ when $\pi^{(k)}$ is an extreme point, and the most informative case is when $\pi^{(k)}$ lies in the relative interior of $\operatorname{OTP}_{c}(\mu^{(k)},\nu^{(k)})$. In particular, if the OT plan $\pi^{(k)}$ is computed with the simplex algorithm, then it is an extreme point and (RI) is least informative.
Remark 2.2 (Computation of (RI)).
Note that (RI) requires explicit knowledge of the extreme points $\operatorname{EP}(\mu,\nu)$. Given a pair of marginals $(\mu,\nu)$, the polytope $\Pi(\mu,\nu)$ is determined via a number of linear equations. Calculating the extreme points from this is called the vertex enumeration problem. There are various algorithms available for this, see e.g. [Avis1992, Bremner1998, Dyer1977] and references therein.
As an instance of an LP, the OT problem has the primal (1) and dual formulation (2). By general LP theory, their solutions are linked in the following way [Luenberger2008, Section 4.4]:
Lemma 2.3 (Primal-dual optimality criterion).
Let $(\mu,\nu)\in\Delta_{N}\times\Delta_{M}$ be a pair of marginals and $c\in\mathbb{R}^{N\times M}$ a cost matrix. A transport plan $\pi\in\Pi(\mu,\nu)$ and potentials $(f,g)\in\Phi_{c}$ are optimal if and only if for all $(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket$ it holds that $\pi_{i,j}>0$ implies $c_{i,j}=f_{i}+g_{j}$.
2.3 links identifiability of the cost matrix $c$ to the support of an OT plan $\pi$, denoted as
$\operatorname{supp}\pi\coloneqq\{(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket:\pi_{i,j}>0\},$
and knowledge of the optimal potentials $f$ and $g$. Upon defining the outer sum
$f\oplus g\coloneqq[f_{i}+g_{j}]_{i,j=1}^{N,M},$
the condition in 2.3 can be stated as $f\oplus g=c$ on $\operatorname{supp}\pi$.
We say that two cost matrices $c,\,c^{\prime}\in\mathbb{R}^{N\times M}$ are shift equivalent if
$c\sim_{\oplus}c^{\prime}\quad\;\vcentcolon\Longleftrightarrow\;\quad c-c^{\prime}=a\oplus b\quad\text{for some vectors }a\in\mathbb{R}^{N}\,,b\in\mathbb{R}^{M}\,.$ (4)
Note that this defines an equivalence relation on $\mathbb{R}^{N\times M}$. Denote with $\operatorname{SC}^{N\times M}\coloneqq\mathbb{R}^{N\times M}/{\sim_{\oplus}}$ the corresponding quotient space. When we say that $c$ is identifiable in $\operatorname{SC}^{N\times M}$, we mean that $c$ is identifiable in $\mathbb{R}^{N\times M}$ up to shift equivalence. This definition is motivated by the fact that for any pair of marginals $(\mu,\nu)\in\Delta_{N}\times\Delta_{M}$ it holds for $c\sim_{\oplus}c^{\prime}$ that $\operatorname{OTP}_{c}(\mu,\nu)=\operatorname{OTP}_{c^{\prime}}(\mu,\nu)$, i.e., the OT w.r.t. shift equivalent cost matrices has the same OT plans. Therefore, for identifiability of the cost matrix $c$ it is crucial to know the total costs i. In particular, when only observing the OT plans ii, we cannot achieve more than identifiability in $\operatorname{SC}^{N\times M}$. We note that the linear space of shifts of the form $a\oplus b$ has dimension $N+M-1$.
Remark 2.4 (Generalization to Linear Programs).
The considerations in this section are a particular case of the more general linear program [Bertsimas1997, Luenberger2008] with fixed (and known) constraint matrix $A\in\mathbb{R}^{E\times D}$: For $\bm{c}\in\mathbb{R}^{D}$, $\bm{b}\in\mathbb{R}^{E}$ consider the primal problem
$\operatorname{LP}_{\bm{c}}(\bm{b})\coloneqq\min_{\bm{x}\in P(\bm{b})}\langle{\bm{c}},{\bm{x}}\rangle\,,$
where the polytope $P(\bm{b})\coloneqq\{\bm{x}\in\mathbb{R}^{D}:A\bm{x}=\bm{b},\,\bm{x}\geq 0\}$ is bounded and non-empty. The corresponding dual problem reads
$\operatorname{LP}_{\bm{c}}(\bm{b})=\max_{\bm{y}\in P^{*}(\bm{c})}\langle{\bm{b}},{\bm{y}}\rangle\,,$
where $P^{*}(\bm{c})\coloneqq\{\bm{y}\in\mathbb{R}^{E}:A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{y}\leq\bm{c}\}$.
Here, we observe $K$ vectors $\bm{b}^{(k)}\in\mathbb{R}^{E}$, $k\in\llbracket K\rrbracket$, with corresponding primal and dual solutions $\bm{x}^{(k)}\in P_{A}(\bm{b^{(k)}})$ and $\bm{y}^{(k)}\in P^{*}(\bm{c})$ as well as the optimal values $\alpha^{(k)}\in\mathbb{R}$, i.e.,
$\alpha^{(k)}=\operatorname{LP}_{\bm{c}}(\bm{b}^{(k)})=\langle{\bm{c}},{\bm{x}^{(k)}}\rangle=\langle{\bm{b}^{(k)}},{\bm{y}^{(k)}}\rangle\,.$
From (a subset of) this information we want to uniquely determine the underlying $\bm{c}$.
Strong duality yields the analogue to 2.3: $\bm{x}\in P(\bm{b})$ and $\bm{y}\in P^{*}(\bm{c})$ are optimal if and only if $\langle{\bm{c}},{\bm{x}}\rangle=\langle{\bm{b}},{\bm{y}}\rangle$, or equivalently if $\bm{x}_{i}>0$ for some $i\in\llbracket D\rrbracket$ implies that $(A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{y})_{i}=\bm{c}_{i}$.
The more general formulation of shift equivalence is as follows: Write $\bm{c}\sim_{A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}\bm{c^{\prime}}$ if $\bm{c}-\bm{c^{\prime}}=A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{y}$ for some $\bm{y}\in\mathbb{R}^{E}$. As it holds for all $\bm{x}\in P(\bm{b})$ that
$\langle{\bm{c}+A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{y}},{\bm{x}}\rangle=\langle{\bm{c}},{\bm{x}}\rangle+\langle{\bm{y}},{A\bm{x}}\rangle=\langle{\bm{c}},{\bm{x}}\rangle+\langle{\bm{y}},{\bm{b}}\rangle\,,$
we see that the shift of $\bm{c}$ by $A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{y}$ does not change the set of minimizers of the linear program. Note that these shifts lie in the image of the linear operator $A^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$.
Finally, we introduce the (to OT specific) concept of Monge matrices (for an overview, see [burkard1996perspectives]). A cost matrix $c\in\mathbb{R}^{N\times M}$ satisfies the so-called Monge property if it holds that
$c_{i,j}+c_{r,s}\leq c_{i,s}+c_{r,j}\qquad\text{for all }1\leq i<r\leq N,\,1\leq j<s\leq M\,.$ (5)
Example 2.5 (Monge matrices).
Let $h:\mathbb{R}\to\mathbb{R}$ be convex and $\{x_{1},\ldots,x_{N}\},\,\{y_{1},\ldots,y_{M}\}$ be increasingly sorted sets of points in $\mathbb{R}$. Then, the induced cost matrix with $c_{i,j}=h(x_{i}-y_{j})$ for all $(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket$ satisfies the Monge property. This includes cost matrices of the form $c(x_{i},y_{j})=\lvert x_{i}-y_{j}\rvert^{p}$ for $p\geq 1$.
When the cost matrix $c$ is Monge, the north-west-corner rule always yields an OT plan [Hoffman1963]. Hence, observing $K$ marginals, we can always construct OT plans $\pi^{(k)}$, $k\in\llbracket K\rrbracket$. To this end, denote the cumulative sums of $\mu^{(k)}$ with
$A^{(k)}_{i}\coloneqq\sum_{r=1}^{i}\mu^{(k)}_{r}\,,\qquad i\in\{0\}\cup\llbracket N\rrbracket\,,$ (6)
and define $\{B^{(k)}_{j}\}_{j=0}^{M}$ similarly as the cumulative sums of $\nu^{(k)}$. Then, an OT plan $\pi^{(k)}$ between $\mu^{(k)}$ and $\nu^{(k)}$ w.r.t. a Monge matrix $c$ is given by the monotone plan
$\pi^{(k)}_{i,j}=\lambda_{1}\{(A_{i-1}^{(k)},A_{i}^{(k)}]\cap(B_{j-1}^{(k)},B_{j}^{(k)}]\}\,,\qquad i\in\llbracket N\rrbracket,\,j\in\llbracket M\rrbracket\,,$ (7)
where $\lambda_{1}$ is the one-dimensional Lebesgue measure. As this does not directly depend on $c$, only on $c$ being Monge, we again see that knowledge of the total cost i is crucial for identifiability.
2.2 Identifiability from Total Costs Alone
First, we consider the most difficult case of only observing the total costs i,
$\alpha^{(k)}=\mathcal{T}_{c}(\mu^{(k)},\nu^{(k)})\,,\qquad k\in\llbracket K\rrbracket\,.$ (8)
Here, we only know that for each $k\in\llbracket K\rrbracket$ there exists at least one extreme point $u^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ such that $\langle{c},{u^{(k)}}\rangle=\alpha^{(k)}$ and any other $v^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ must satisfy $\langle{c},{v^{(k)}}\rangle\geq\alpha^{(k)}$. This leads to a multitude of LPs whose solutions, if existent, all give a valid cost matrix.
Theorem 2.6.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i. Then, $c$ is identifiable in $\mathbb{R}^{N\times M}$ if and only if for all possible combinations of $u^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$, $k\in\llbracket K\rrbracket$, the system
$\langle{c},{v^{(k)}}\rangle$ $\geq\alpha^{(k)},$ $v^{(k)}$ $\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})\setminus\{u^{(k)}\},$ (9)
$\langle{c},{u^{(k)}}\rangle$ $=\alpha^{(k)},$ $k$ $\in\llbracket K\rrbracket,$
is either inconsistent or the solution is unique and independent of the choice. Note that if a solution exists (that is allowed to depend on the choice), then we have existence of the underlying cost matrix $c$.
Proof.
A cost matrix $c$ is consistent with the observed OT information if and only if it satisfies the above system (9) for at least one combination. ∎
Remark 2.7 (Additional linear constraints on $c$).
Each system (9) in Theorem 2.6 is an LP. This formulation is quite flexible as it allows us to place additional linear (in-)equality constraints on the cost matrix $c$. However, the LP might not be proper in the sense that the feasibility set can be unbounded. In general, this can be dealt with by placing a boundedness restriction on the set of cost matrices $c$, e.g., $0\leq c\leq C_{0}$ for some constant $C_{0}>0$.
Remark 2.8 (Set of consistent cost matrices).
As each system (9) in Theorem 2.6 is an LP, it follows that the set of consistent cost matrices is equal to a union of at most $K$ (possibly unbounded) polytopes.
Remark 2.9 (Reformulation as Mixed Integer Linear Program).
Note that the formulation involving all possible combinations of optimal extreme points in Theorem 2.13 can be reformulated as follows: The cost matrix $c$ is identifiable in $\mathbb{R}^{N\times M}$ if and only if the system
$\langle{c},{v^{(k)}}\rangle\geq\alpha^{(k)}\,,\qquad v^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)}),\qquad k\in\llbracket K\rrbracket,$
where equality is attained for at least one $v^{(k)}$ for each $k\in\llbracket K\rrbracket$, has a unique solution.
Assuming that $c$ is bounded by $C_{0}>0$, this can be formulated as a mixed integer linear program (MILP), i.e., an LP where some variables are constrained to be integer. To this end, enumerate for all $k\in\llbracket K\rrbracket$ the extreme points $\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ as $\{v^{(1,k)},\ldots,v^{(R_{k},k)}\}$ and introduce for $j\in\llbracket R_{k}\rrbracket$, $k\in\llbracket K\rrbracket$ the binary variables $e^{(j,k)}\in\{0,1\}$ and the slack variables $z^{(j,k)}\in\mathbb{R}$. Then, the MILP reads
$\sum_{r=1}^{R_{k}}e^{(r,k)}$ $\geq 1,$ $0\leq z^{(j,k)}$ $\leq[C_{0}-\alpha^{(k)}](1-e^{(j,k)}),$
$\langle{c},{v^{(j,k)}}\rangle-z^{(j,k)}$ $=\alpha^{(k)},$ $j$ $\in\llbracket R_{k}\rrbracket,\;k\in\llbracket K\rrbracket.$
Note that the free variables are $c$ and $e^{(j,k)},\,v^{(j,k)}$ with $j\in\llbracket R_{k}\rrbracket$, $k\in\llbracket K\rrbracket$. In particular, $c$ is identifiable in $[0,C_{0}]^{N\times M}$ if and only if the MILP has a unique solution in $c$ (not the full set of free variables). Since solving MILPs is NP-hard, in general, this suggests that inferring the cost matrix from marginals and total costs only is a difficult and computationally expensive problem. The situation simplifies significantly if the OT plans are observed, see Section 3. Furthermore, note that in the current formulation, any $e^{(k,r)}=1$ can be changed to 00 if $\sum_{j=1}^{R_{k}}e^{(j,k)}\geq 2$, without changing the solution in $c$. Hence, uniqueness of the full solution of the MILP is stronger than identifiability of $c$. In fact, the MILP can only have a unique solution when the underlying optimal faces each consist of a single point.
Remark 2.10 (Simpler sufficient condition).
By dropping the inequalities in the system (9) given in Theorem 2.6, we get a sufficient condition for identifiability that is easier to check as it only involves linear equations. Namely, $c$ is identifiable in $\mathbb{R}^{N\times M}$ if for all possible combinations of potentially optimal $u^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$, $k\in\llbracket K\rrbracket$, the system
$\langle{c},{u^{(k)}}\rangle=\alpha^{(k)},\qquad k\in\llbracket K\rrbracket,$ (10)
is either inconsistent or the solution is unique and independent of the choice. Note that here it is crucial that we assume that the underlying cost matrix $c$ exists, as the above system (in contrast to the one in Theorem 2.6) does not ensure consistency, see also 2.27.
We now give an example where Theorem 2.6 provides identifiability but not the conditions given in 2.10.
Example 2.11.
Suppose that for $N=M=2$ we observe the following OT information consisting of $K=4$ pairs of marginals and corresponding total costs
$\mu^{(1)}$ $=(3/4,1/4),$ $\mu^{(2)}$ $=(3/7,4/7),$ $\mu^{(3)}$ $=(4/5,1/5),$ $\mu^{(4)}$ $=(3/7,4/7),$
$\nu^{(1)}$ $=(5/8,3/8),$ $\nu^{(2)}$ $=(1/5,4/5),$ $\nu^{(3)}$ $=(1/2,1/2),$ $\nu^{(4)}$ $=(1/2,1/2),$
$\alpha^{(1)}$ $=7/4,$ $\alpha^{(2)}$ $=1,$ $\alpha^{(3)}$ $=1,$ $\alpha^{(4)}$ $=1.$
Then, it can be checked that Theorem 2.6 yields identifiability of the underlying cost matrix $c$ in $\mathbb{R}^{N\times M}$, and it is given by
$c=\begin{bmatrix}9/2&-2\\ 13/4&13/4\end{bmatrix}.$
To be more precise, only the following combination of extreme points yields a solution to the system (9):
$u^{(1)}=\begin{bmatrix}3/8&3/8\\ 1/4&0\end{bmatrix},\quad u^{(2)}=\begin{bmatrix}0&3/7\\ 1/5&13/35\end{bmatrix},\quad u^{(3)}=\begin{bmatrix}3/10&1/2\\ 1/5&0\end{bmatrix},\quad u^{(4)}=\begin{bmatrix}0&3/7\\ 1/2&1/14\end{bmatrix}.$
All other combinations of extreme points lead to an inconsistent system (9). Notably, when dropping the inequality constraints, each corresponding system (10) has an unique solution that is not independent of the combination. Hence, in this example 2.10 cannot be used to obtain identifiability.
If we know that the cost matrix satisfies the Monge property (5), we already know one extreme point that is optimal. Hence, in this case identifiability reduces to the unique solvability of a single LP.
Corollary 2.12.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i. For each $k\in\llbracket K\rrbracket$ denote with $\pi^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ the monotone transport plan (7). Then, $c$ is identifiable in the class of Monge matrices if and only if the system (9) with $u^{(k)}=\pi^{(k)}$, $k\in\llbracket K\rrbracket$, has a unique solution. Note that if a solution exists that satisfies the Monge property, then we have existence of the underlying cost matrix $c$.
Note that the Monge property (5) is a linear inequality constraint on the cost matrix $c$. Hence, the Monge matrix constraint can be explicitly added to the system (9) of 2.12 without changing the LP structure.
2.3 Identifiability from Optimal Potentials Alone
Next, we consider the case where we only know the optimal potentials iii. Note that, in such a scenario, the total costs i are determined by the relations
$\alpha^{(k)}=\mathcal{T}_{c}(\mu^{(k)},\nu^{(k)})=\langle{f^{(k)}},{\nu^{(k)}}\rangle+\langle{g^{(k)}},{\nu^{(k)}}\rangle\,,\qquad k\in\llbracket K\rrbracket\,.$ (11)
We proceed as in Theorem 2.6 and consider all combinations of possible optimal extreme points. Using the primal-dual optimality criterion (Theorem 2.17), we get a more restrictive system.
Theorem 2.13.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i and optimal potentials iii. Then, $c$ is identifiable in $\mathbb{R}^{N\times M}$ if and only if for all possible combinations of $u^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$, $k\in\llbracket K\rrbracket$, the system
$\langle{c},{v^{(k)}}\rangle$ $\geq\alpha^{(k)},$ $\langle{c},{u^{(k)}}\rangle$ $=\alpha^{(k)},$ (12)
$c$ $\geq f^{(k)}\oplus g^{(k)}\text{ with equality on }\operatorname{supp}u^{(k)},$ $k$ $\in\llbracket K\rrbracket,$
$v^{(k)}$ $\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})\setminus\{u^{(k)}\},$
is either inconsistent or the solution is unique and independent of the choice. Note that if a solution exists (that is allowed to depend on the choice), then we have existence of the underlying cost matrix $c$.
For the class of Monge matrices, we can again use the structural knowledge of the OT to arrive the following more specialized version of Theorem 2.13.
Corollary 2.14.
Suppose that we observe $K$ pairs of marginals and the corresponding total costs i and optimal potentials iii. For each $k\in\llbracket K\rrbracket$ denote with $\pi^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})$ the monotone transport plan (7). Then, the cost matrix $c$ is identifiable in the class of Monge matrices if and only if the system (12) with $u^{(k)}=\pi^{(k)}$, $k\in\llbracket K\rrbracket$, has a unique solution. Note that if a solution exists that satisfies the Monge property, then we have existence of the underlying cost matrix.
Only using the primal-dual optimality criterion (2.3), we obtain the following easily verifiable sufficient condition for the identifiability of Monge matrices. Recall that this encompasses all the matrices given in 2.5.
Corollary 2.15.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i and optimal potentials iii. Recall the definition of $A^{(k)}_{i}$ and $B^{(k)}_{j}$ in (6). Then, $c$ is identifiable in the class of Monge matrices if for all $i\in\llbracket N\rrbracket$, $j\in\llbracket M\rrbracket$ there exists a $k\in\llbracket K\rrbracket$ such that
$\max(A_{i-1}^{(k)},B_{j-1}^{(k)})<\min(A_{i}^{(k)},B_{j}^{(k)})\,.$
Next, we give an example where 2.14, but not 2.15, yields identifiability.
Example 2.16.
Suppose that for $N=M=2$ we observe the following OT information consisting of $K=3$ pairs of marginals and corresponding total costs and optimal potentials
$\mu^{(1)}$ $=(1/2,1/2),$ $\mu^{(2)}$ $=(1/3,2/3),$ $\mu^{(3)}$ $=(1,0),$
$\nu^{(1)}$ $=(1,0),$ $\nu^{(2)}$ $=(1,0),$ $\nu^{(3)}$ $=(2/5,3/5),$
$\alpha^{(1)}$ $=1,$ $\alpha^{(2)}$ $=2/3,$ $\alpha^{(3)}$ $=5/3,$
$f^{(1)}$ $=(0,-2),$ $f^{(2)}$ $=(0,-2),$ $f^{(3)}$ $=(2,0),$
$g^{(1)}$ $=(2,0),$ $g^{(2)}$ $=(2,0),$ $g^{(3)}$ $=(0,-5/9).$
Then, 2.14 yields that the following underlying cost matrix
$c=\begin{bmatrix}2&13/9\\ 0&-5/9\end{bmatrix}$
is identifiable in the class of Monge matrices. However, the given OT information is not enough to obtain identifiability via 2.15.
2.4 Identifiability from Optimal Transport Plans Alone
This subsection is devoted to the identifiability of the cost matrix when only the OT plans ii are observed. This is quite different from the other settings as due to the shift invariance (4) we cannot achieve true identifiability. Nevertheless, we will be able to identify the underlying cost matrix to some extent.
Figure 1: The polytope $\Pi(\mu,\nu)$ (gray-filled) with optimality set $\operatorname{OTP}_{c}(\mu,\nu)$ (dotted line): The observation of a point in the relative interior of the optimality set provides more information than the knowledge of an extreme point. Middle figure: Only one possible tangent plane (dashed line) at the point $\pi$. Right figure: A whole bundle of hyperplanes (dashed lines). Therefore, in the figure in the middle, we have more information to identify the cost matrix.
Observing the optimal plans ii, we know by (RI) that the smallest face that $\pi^{(k)}$ lies on must be contained in $\operatorname{OTP}_{c}(\mu^{(k)},\nu^{(k)})$. This yields further information about the tangent hyperplane $\pi^{(k)}+\{x\in\mathbb{R}^{N\times M}:\langle{x},{c}\rangle=0\}$, see Figure 1. Depending on the size of $\operatorname{span}\{\pi^{(k)}-u^{(\ell,k)}:k\in\llbracket K\rrbracket,\,\ell\in\llbracket L_{k}\rrbracket\}$, we will see that the cost matrix may be identifiable in a suitable sense. Recall that $\operatorname{SC}^{N\times M}$ denotes the quotient space of cost matrices up to shift equivalence (4). In the following theorem, we show that we can identify $c$ in $\operatorname{SC}^{N\times M}$ up to a span of a number of linearly independent vectors.
Theorem 2.17.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding OT plans ii. Let $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$, be as in (RI). Then, the following holds:
  1. 1.
    Let $S\in\mathbb{N}_{0}$ be such that
    $\dim(\operatorname{span}\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\})=(N-1)(M-1)-S,$
    then, the cost matrix $c$ is identifiable in $\operatorname{SC}^{N\times M}$ up to the span of $S$ linearly independent vectors.
  2. 2.
    If $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$ for all $k\in\llbracket K\rrbracket$, then the converse holds.
Proof.
For $k\in\llbracket K\rrbracket$ it holds by $\pi^{(k)},u^{(1,k)},\ldots,u^{(L_{k},k)}\in\operatorname{OTP}_{c}(\pi^{(k)})$ that
$\langle{c},{\pi^{(k)}}\rangle=\langle{c},{u^{(\ell,k)}}\rangle\qquad\text{for all }\ell\in\llbracket L_{k}\rrbracket\,.$
This implies that
$\operatorname{span}\left\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}\perp c$
and as a consequence,
$c\in\left[\operatorname{span}\left\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}\right]^{\perp}\,.$
By assumption, this orthogonal complement has dimension
$NM-(N-1)(M-1)+S=N+M+S-1.$
We show that $N+M-1$ of these dimensions are spanned by $\{a\oplus b:a\in\mathbb{R}^{N},\;b\in\mathbb{R}^{M}\}$. To this end, we prove that $\left\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}$ and $\{a\oplus b:a\in\mathbb{R}^{N},\;b\in\mathbb{R}^{M}\}$ are linearly independent. For a fixed $k\in\llbracket K\rrbracket$, suppose that there exists scalars $\lambda^{(k)}_{\ell}\in\mathbb{R}$, $\ell\in\llbracket L_{k}\rrbracket$, and $R\in\mathbb{N}$ pairs $(a^{(r)},b^{(r)})\in\mathbb{R}^{N}\times\mathbb{R}^{M}$ with scalars $\tau_{r}\in\mathbb{R}$, $r\in\llbracket R\rrbracket$, such that
$v\coloneqq\sum_{k=1}^{K}\sum_{\ell=1}^{L_{k}}\lambda^{(k)}_{\ell}[\pi^{(k)}-u^{(\ell,k)}]=\sum_{r=1}^{R}\tau_{r}[a^{(r)}\oplus b^{(r)}]\,.$
Then, it follows that
$\lVert v\rVert_{2}^{2}$ $=\left\langle{\sum_{k=1}^{K}\sum_{\ell=1}^{L_{k}}\lambda^{(k)}_{\ell}[\pi^{(k)}-u^{(\ell,k)}]},{\sum_{r=1}^{R}\tau_{r}[a^{(r)}\oplus b^{(r)}]}\right\rangle$
$=\sum_{k=1}^{K}\sum_{\ell=1}^{L_{k}}\sum_{r=1}^{R}\lambda^{(k)}_{\ell}\tau_{r}\langle{\pi^{(k)}-u^{(\ell,k)}},{a^{(r)}\oplus b^{(r)}}\rangle$
$=\sum_{k=1}^{K}\sum_{\ell=1}^{L_{k}}\sum_{r=1}^{R}\lambda^{(k)}_{\ell}\tau_{r}\left[\sum_{i=1}^{N}a_{i}^{(r)}(\mu^{(k)}_{i}-\mu^{(k)}_{i})+\sum_{j=1}^{M}b_{j}^{(r)}(\nu^{(k)}_{j}-\nu^{(k)}_{j})\right]=0\,,$
which implies $v=0$ and a fortiori $\left\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}$ and $\{a\oplus b:a\in\mathbb{R}^{N},\;b\in\mathbb{R}^{M}\}$ are linearly independent. Hence, $c$ is determined in $\operatorname{SC}^{N\times M}$ up to the remaining $S$ dimensions.
For the converse, assume that
$\dim(\operatorname{span}\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\})<(N-1)(M-1)-S\,.$
Then, there exists more than $S$ linearly independent vectors, so that there exist $c^{\prime}$ such that
$c,c^{\prime}\in\left[\operatorname{span}\left\{\pi^{(k)}-u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}\right]^{\perp}\,.$
Then,
$\langle{c},{\pi^{(k)}-u^{(\ell,k)}}\rangle=0=\langle{c^{\prime}},{\pi^{(k)}-u^{(\ell,k)}}\rangle\qquad\text{for all }\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket.$ (13)
Hence,
$\langle{c},{\pi^{(k)}}\rangle=\langle{c},{u^{(\ell,k)}}\rangle\quad{\rm and}\quad\langle{c^{\prime}},{\pi^{(k)}}\rangle=\langle{c^{\prime}},{u^{(\ell,k)}}\rangle\qquad\text{for all }\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket.$
We claim that there exists $\tau(k)>0$ such that for all $0\leq\tau\leq\tau(k)$ it holds that $\pi^{(k)}\in\operatorname{OTP}_{c_{\tau}}(\pi^{(k)})$ where $c_{\tau}=c+\tau c^{\prime}$. We prove it by contradiction. Assume that for all $\tau(k)>0$ there exists $0\leq\tau\leq\tau(k)$ such that $\pi^{(k)}\notin\operatorname{OTP}_{c_{\tau}}(\pi^{(k)})$. Then, we can create a sequence of positive numbers $\tau_{s}>0$ with $\tau_{s}\to 0$ for $s\to\infty$ and a sequence of couplings $\gamma_{s}\in\operatorname{OTP}_{c_{\tau_{s}}}(\pi^{(k)})\cap\operatorname{EP}(\pi^{(k)})$ such that
$\langle{c_{\tau_{s}}},{\pi^{(k)}}\rangle>\langle{c_{\tau_{s}}},{\gamma_{s}}\rangle.$ (14)
We can further assume (subtracting, if necessary, a further subsequence) that $\gamma_{s}\to\gamma\in\Pi(\pi^{(k)})$ as $s\to\infty$. Since the set of extreme points is finite, it follows that $\gamma\in\operatorname{EP}(\pi^{(k)})$ and $\gamma_{s}=\gamma$ for all $s\geq s_{0}$ for some $s_{0}\in\mathbb{N}$. Since, for any $\pi\in\Pi(\pi^{(k)})$ it holds that $\langle{c_{\tau_{s}}},{\gamma_{s}}\rangle\leq\langle{c_{\tau_{s}}},{\pi}\rangle$, we have
$\langle{c_{\tau_{s}}},{\gamma_{s}-\gamma}\rangle+\langle{c_{\tau_{s}}},{\gamma}\rangle\leq\langle{c_{\tau_{s}}},{\pi}\rangle.$
Taking limits, we get
$\langle{c},{\gamma}\rangle\leq\langle{c},{\pi}\rangle,\qquad\text{ for all }\pi\in\Pi(\pi^{(k)})\,,$
so that $\gamma\in\operatorname{OTP}_{c}(\pi^{(k)})\cap\operatorname{EP}(\pi^{(k)})$. As $\operatorname{coh}\left(\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}\right)=\operatorname{OTP}_{c}(\pi^{(k)})$, we can assume without loss of generality that $\gamma=u^{(1,k)}$. We notice that in such a case
$\langle{c},{\pi^{(k)}-u^{(1,k)}}\rangle=0.$
Using (13), we have for all $s\geq s_{0}$ that
$\langle{c_{\tau_{s}}},{\gamma_{s}}\rangle=\langle{c},{u^{(1,k)}}\rangle+\tau_{s}\langle{c^{\prime}},{u^{(1,k)}}\rangle=\langle{c},{\pi^{(k)}}\rangle+\tau_{s}\langle{c^{\prime}},{\pi^{(k)}}\rangle=\langle{c_{\tau_{s}}},{\pi^{(k)}}\rangle\,,$
which contradicts (14). Therefore, the claim holds. Then, for any $k\in\llbracket K\rrbracket$, the observed plan $\pi^{(k)}$ is optimal for any of cost matrix $c_{\tau}=c+\tau\,c^{\prime}$, with $0\leq\tau\leq\min\{\tau(1),\ldots,\tau(K)\}$, so that identifiability cannot hold in $\operatorname{SC}^{N\times M}$ up to the span of $S$ linearly independent vectors. ∎
Remark 2.18 (The case $S=0$).
In Theorem 2.17 with $S=0$, the given dimension of $(N-1)(M-1)$ can be attained. Namely, a single transport polytope can have this affine dimension [Deloera2013]. Note that in this case it follows for the underlying cost matrix that $c\sim_{\oplus}0$.
Remark 2.19 (The case $S=1$).
In Theorem 2.17, if $S=1$ then the cost matrix $c$ is determined up to shift equivalence and multiplication by a scalar, i.e., $c$ can (up to shift equivalence) be written as $c=\tau c^{\prime}$ for $\tau>0$ and some cost matrix $c^{\prime}$ that has the right sign (determined such that all the observed OT plans are optimal for $c^{\prime}$). If we care only about calculating the OT plans, then the true value of $\tau$ does not matter.
Remark 2.20 (Computation of the span containing $c$).
The proof of Theorem 2.17 reveals how the span which contains $c$ (up to shift equivalence) can be calculated from the observed OT plans. In fact, this can be implemented by solving a linear system of equations. Let
$E^{(k)}\coloneqq\begin{pmatrix}[\pi^{(k)}-u^{(1,k)}]^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\\ \vdots\\ [\pi^{(k)}-u^{(L_{k},k)}]^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\\ \end{pmatrix}\,,\qquad E_{K}\coloneqq\begin{pmatrix}E^{(1)}\\ \vdots\\ E^{(K)}\end{pmatrix}\,.$
Then, $c$ must satisfy
$E_{K}c=0\,.$
2.5 Identifiability from Total Costs and Optimal Transport Plans
Now, we consider the case where we only observe the total costs i and OT plans ii, i.e.,
$\alpha^{(k)}=\mathcal{T}_{c}(\mu^{(k)},\nu^{(k)})=\langle{c},{\pi^{(k)}}\rangle\,,\quad k\in\llbracket K\rrbracket\,.$ (15)
This setting is much more simple than that in Theorem 2.6 or Theorem 2.13, as we directly observe OT plans and therefore do not have to consider all possible ones. Recall that by (RI) the observed OT plans $\pi^{(k)}$ potentially contain more information depending on where they lie in the corresponding polytope $\Pi(\pi^{(k)})$.
Theorem 2.21.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i and OT plans ii. Let $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$, be as in (RI). Then, the underlying cost matrix $c$ is identifiable in $\mathbb{R}^{N\times M}$ if and only if the system
$\langle{c},{v^{(k)}}\rangle$ $\geq\alpha^{(k)},$ $v^{(k)}$ $\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})\setminus\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\},$
$\langle{c},{u^{(\ell,k)}}\rangle$ $=\alpha^{(k)},$$\ell$ $\in\llbracket L_{k}\rrbracket,\;k\in\llbracket K\rrbracket,$
has a unique solution. Note that if a solution exists, then we have existence of the underlying cost matrix.
The system in Theorem 2.21 is an LP. In particular, identifiability is thus linked to the uniqueness of solutions of LPs [Mangasarian1979, Appa2002]. Dropping the inequalities in the system of Theorem 2.21, we get a sufficient condition for identifiability in terms of a linear equation:
$\langle{c},{u^{(\ell,k)}}\rangle=\alpha^{(k)},\qquad\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket.$ (16)
From this we can get a simple identifiability criterion in terms of the OT plans.
Theorem 2.22.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i and OT plans ii. Let $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$, be as in (RI). Then, the following holds:
  1. a)
    If the system $\{u^{(\ell,k)}:k\in\llbracket K\rrbracket,\,\ell\in\llbracket L_{k}\rrbracket\}$ contains $NM$ linearly independent vectors, then the cost matrix $c$ is identifiable in $\mathbb{R}^{N\times M}$.
  2. b)
    If $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$ for all $k\in\llbracket K\rrbracket$, then the converse holds.
Proof.
If the system contains $NM$ linearly independent vectors, then the solution $c$ to (16) is fully determined.
For the converse, suppose that the system $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\}$ contains less than $NM$ linearly independent vectors. In particular, there exists $c^{\prime}\neq 0$ in the kernel of the linear equations (16). Define for $\tau>0$ the cost matrix $c_{\tau}\coloneqq c+\tau c^{\prime}$. We claim that for all $k\in\llbracket K\rrbracket$ there exists a $\tau(k)>0$ such that for all $0<\tau\leq\tau(k)$ it holds that $\pi^{(k)}$ is optimal w.r.t. the cost matrix $c_{\tau}$. Fix $k\in\llbracket K\rrbracket$ and assume that this is not the case, i.e., for all $\tau(k)>0$ there exists a $0<\tau\leq\tau(k)$ such that $\langle{\pi^{(k)}},{c_{\tau}}\rangle=\alpha^{(k)}>\mathcal{T}_{c_{\tau}}(\mu^{(k)},\nu^{(k)})$. Hence, we can construct sequences $\tau_{s}>0$ and $\gamma_{s}\in\operatorname{OTP}_{c_{\tau_{s}}}(\pi^{(k)})\cap\operatorname{EP}(\pi^{(k)})$ such that $\alpha^{(k)}>\langle{c_{\tau_{s}}},{\gamma_{s}}\rangle$ and $\tau_{s}\to 0$ for $s\to\infty$. As in the proof of Theorem 2.17 it follows that there exist $\gamma\in\operatorname{OTP}_{c}(\pi^{(k)})$ and $s_{0}\in\mathbb{N}$ such that $\gamma_{s}=\gamma$ for all $s\geq s_{0}$. If $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$, then $\gamma=u^{(\ell,k)}$ for some $\ell\in\llbracket L_{k}\rrbracket$ and thus $\langle{c^{\prime}},{\gamma}\rangle=0$. Hence, $\langle{c_{\tau}},{\gamma_{s}}\rangle=\langle{c_{\tau}},{\gamma}\rangle=\langle{c},{\gamma}\rangle=\alpha^{k}$, a contradiction. Therefore, for any $k\in\llbracket K\rrbracket$, the observed plan $\pi^{(k)}$ is optimal for the cost matrix $c_{\tau}=c+\tau c^{\prime}$, with $0\leq\tau\leq\min\{\tau(1),\ldots,\tau(K)\}$, so that $c$ is not identifiable. This concludes the proof. ∎
Remark 2.23 (Repeated pairs of marginals).
In Theorem 2.21 and Theorem 2.22 we assume that the $K$ observed pairs of marginals are distinct. If they are not, say $(\mu^{(1)},\nu^{(1)})=\ldots=(\mu^{(R)},\nu^{(R)})$, then we can set the OT plan of this pair of marginals to $\frac{1}{R}\sum_{r=1}^{R}\pi^{(r)}$ and both theorems remain valid. Note that this joins the information of the individual OT plans $\pi^{(1)},\ldots,\pi^{(R)}$ in the sense that the resulting face given in (RI) is maximal in terms of the available information.
Remark 2.24 (Shift equivalence and total costs).
Although shift equivalent (4) cost matrices induce the same OT plans, the total cost may differ. Hence, in the setting where we also observe the total costs, we can achieve identifiability in $\mathbb{R}^{N\times M}$ and not just $\operatorname{SC}^{N\times M}$.
However, this setting can still be influenced by a stronger version of shift equivalency given a fixed number of marginals: Assume that there exist $a\in\mathbb{R}^{N}$ and $b\in\mathbb{R}^{M}$ such that for all $k\in\llbracket K\rrbracket$ it holds that
$\langle{a},{\mu^{(k)}}\rangle+\langle{b},{\nu^{(k)}}\rangle=0\,.$
Then, the shift $a\oplus b$ does not change the total cost for these specific $K$ marginals. Hence, when one of $a$ or $b$ is non-zero, then we have another cost matrix $c+a\otimes b\neq c$ that has the same OT plans and total costs for the given $K$ pairs of marginals. The condition above can be written as
$Aa+Bb=0,$
where $A\in\mathbb{R}^{K\times N}$ and $B\in\mathbb{R}^{K\times M}$ have the rows $\mu^{(1)},\ldots,\mu^{(K)}$ and $\nu^{(1)},\ldots,\nu^{(K)}$, respectively. This has a non-trivial solution $(a,b)\neq 0$ if and only if
$\ker A\neq\{0\}\quad\text{or}\quad\ker B\neq\{0\}\quad\text{or}\quad\operatorname{im}A\cap\operatorname{im}(-B)\neq\{0\}\,.$
We now illustrate Theorem 2.21 and Theorem 2.22 with some examples. The first one shows that there always exist OT plans and marginals such that the underlying cost matrix $c$ is identifiable.
Example 2.25.
Let $K=NM$ and suppose that we observe the OT plans
$\pi^{(i+[j-1]M)}=\delta_{i}\otimes\delta_{j}\,,\quad i\in\llbracket N\rrbracket,\,j\in\llbracket M\rrbracket\,,$
with corresponding total costs $\alpha^{(k)}$, $k\in\llbracket K\rrbracket$. Here, $\delta_{i}$ and $\delta_{j}$ denote the Dirac measures at $i$ and $j$, respectively. Then, each $\bm{\pi}^{(k)}$ is equal to the $k$-th standard basis vector $\bm{e}_{k}$ and $c$ is therefore identifiable with $c_{i,j}=\alpha^{(i+[j-1]M)}$.
Example 2.26.
We give an example that shows that for Theorem 2.22 the additional assumption of $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$ for all $k\in\llbracket K\rrbracket$ for the converse direction is crucial. Suppose that for $N=M=2$ we observe the $K=3$ OT plans
$\pi^{(1)}=\begin{bmatrix}3/20&1/4\\ 3/5&0\end{bmatrix}\,,\quad\pi^{(2)}=\begin{bmatrix}0&1/2\\ 1/2&0\end{bmatrix}\,,\quad\pi^{(3)}=\begin{bmatrix}1/2&1/10\\ 0&2/5\end{bmatrix}\,,$
with corresponding total costs $\alpha^{(1)}=1/3$ and $\alpha^{(2)}=\alpha^{(3)}=1$. As $K<NM$, we cannot use Theorem 2.22 to obtain identifiability of the underlying cost matrix in $\mathbb{R}^{N\times M}$. However, in this case Theorem 2.21 yields identifiability and the underlying cost matrix is given by
$c=\begin{bmatrix}-1/3&7/3\\ -1/3&7/3\end{bmatrix}\,.$
Note for Theorem 2.22 it is crucial that we know that the underlying cost matrix $c$ exists. Else, it could happen that the observed OT plans are not optimal for the unique $c$ solving the system of linear equations (16).
Example 2.27.
Consider the $K=4$ linearly independent transport plans
$\pi^{(1)}=\begin{bmatrix}0&1/4\\ 1/2&1/4\end{bmatrix},\;\pi^{(2)}=\begin{bmatrix}1/3&0\\ 1/3&1/3\end{bmatrix},\;\pi^{(3)}=\begin{bmatrix}1/3&1/3\\ 0&1/3\end{bmatrix},\;\pi^{(4)}=\begin{bmatrix}4/11&3/11\\ 4/11&0\end{bmatrix}$
with total costs $\alpha^{(1)}=\alpha^{(2)}=1$ and $\alpha^{(3)}=\alpha^{(4)}=2$. Then, it follows that the unique solution to (16) is given by
$c=\begin{bmatrix}8/3&10/3\\ 1/3&0\end{bmatrix}\,.$
However, the total costs for the marginals of the above transport plans w.r.t. $c$ are given by $3/4$, $1$, $2$ and $19/11$, respectively.
Lastly, we note that in Theorem 2.21 it is not necessary to enforce the constraint $\langle{c},{v^{(k)}}\rangle\geq\alpha^{(k)}$ for all $v^{(k)}\in\operatorname{EP}(\mu^{(k)},\nu^{(k)})\setminus\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$. Indeed, according to the next lemma it is enough to consider the extreme points of the biggest faces containing $\pi^{(k)}$, $k\in\llbracket K\rrbracket$.
Lemma 2.28.
Let $c\in\mathbb{R}^{N\times M}$ be a cost function. Suppose that we observe marginals $(\mu,\nu)\in\Delta_{N}\times\Delta_{M}$ with optimal face $F\subseteq\operatorname{OTP}_{c}(\mu,\nu)$ and total cost $\alpha=\mathcal{T}_{c}(\mu,\nu)$. Denote with $G_{1},\ldots,G_{R_{F}}$ the sets containing the extreme points of the highest dimensional faces of $\Pi(\mu,\nu)$ (excepting $\Pi(\mu,\nu)$ itself) that contain $F$ and define the set
$G\coloneqq\begin{cases}\bigcup_{r=1}^{R_{F}}G_{r}\setminus F&R_{F}\neq 1\,,\\ \{w\}&R_{F}=1\,,\text{ for any }w\in\operatorname{EP}(\mu,\nu)\setminus F\,.\end{cases}$
Then, the set of cost functions $c$ that is consistent with the observed OT information is given by the solutions to the system
$\langle{c},{u}\rangle=\alpha,\qquad\langle{c},{v}\rangle\geq\alpha,\qquad u\in F,\qquad v\in G.$
Proof.
Let $c$ be any solution to the above system, then we have to show that $\alpha=\mathcal{T}_{c}(\mu,\nu)$. First, we do the following reduction to the dimension of $\Pi(\mu,\nu)$: Let $B\coloneqq\operatorname{aff}\{\Pi(\mu,\nu)\}$ be the minimal affine space that $\Pi(\mu,\nu)$ lies in (i.e. its affine hull) and denote with $b\coloneqq\dim B$ its dimension. Consider the hyperplane $H\coloneqq\{x\in\mathbb{R}^{NM}:\langle{c},{x}\rangle=\alpha\}$, then if $B\cap H=B$ we are done, else $B\cap H\subsetneq B$ which implies that $H^{\prime}\coloneqq B\cap H$ is a $b$-dimensional hyperplane in $B$ with half-spaces $H^{\prime}_{\geq}\coloneqq\{x\in B:\langle{c},{x}\rangle\geq\alpha\}$ and $H^{\prime}_{\leq}\coloneqq\{x\in B:\langle{c},{x}\rangle\leq\alpha\}$. In particular, the hyperplane $H^{\prime}$ is determined by $b-1$ points that don’t lie on a lower-dimensional hyperplane. Now, we consider the two cases $R_{F}=1$ and $R_{F}>1$ separately.
If $R_{F}=1$, this means that $F=G_{1}$ is maximum dimensional, i.e., it has dimension $b-1$, and $H^{\prime}$ is thus determined by $F$. Again using that $F$ is a maximum dimensional face, it follows that either $\Pi(\mu,\nu)\subseteq H^{\prime}_{\geq}$ or $\Pi(\mu,\nu)\subseteq H^{\prime}_{\leq}$. Due to the constraint $\langle{c},{w}\rangle\geq\alpha$ for one $w\in\operatorname{EP}(\mu,\nu)/F$, we must have $\Pi(\mu,\nu)\subseteq H^{\prime}_{\geq}$ and thus $\alpha=\mathcal{T}_{c}(\mu,\nu)$.
If $R_{F}>1$, we first consider the case that $\langle{c},{v}\rangle=\alpha$ for all $v\in G_{i}$ for some $i\in\llbracket R_{F}\rrbracket$. As $G_{1},\ldots,G_{R_{F}}$ are maximum dimensional, it follows that there exists exactly one $i\in\llbracket R_{F}\rrbracket$ such that $\langle{c},{v}\rangle=\alpha$ for all $v\in G_{i}$. Hence, we can argue as in the $R_{F}=1$ case that $\alpha=\mathcal{T}_{c}(\mu,\nu)$. The other case is $\langle{c},{v}\rangle>\alpha$ for some $v\in G_{i}$ and all $i\in\llbracket R_{F}\rrbracket$. Suppose that $c$ is not consistent with the observed OT information, i.e., there exists $z\in\operatorname{EP}(\mu,\nu)$ with $\langle{c},{z}\rangle<\alpha$. Since by construction $\langle{c},{v}\rangle\geq\alpha$ for all $v\in\bigcup_{i=1}^{R_{F}}G_{i}$, note that $z\notin\bigcup_{i=1}^{R_{F}}G_{i}$. Fix a point $y$ in the relative interior of $F$ and consider the segment $[y,z]\coloneqq\{\lambda y+(1-\lambda)z:0\leq\lambda\leq 1\}\subseteq\Pi(\mu,\nu)$ between $y$ and $z$. As $F$ has dimension strictly less than $b-1$ (else we would have $R_{F}=1$), there exists a $2$-dimensional affine space $E$ that contains $[y,z]$ and is orthogonal to $F$. Then, $\Pi(\mu,\nu)\cap E$ is a $2$-dimensional polytope and we are in the setting of Figure 2. Now, note that the segments $[y,w_{1}]$ and $[y,w_{2}]$ must belong to some $G_{i}$ as they are of maximal dimension. Hence, Figure 2 shows a contradiction: It holds that $\langle{c},{x}\rangle<\alpha$ for all $x\in[y,z]\setminus\{y,z\}$ (the dotted line), but $\langle{c},{x}\rangle\geq\alpha$ for all $x\in\operatorname{coh}\{y,w_{1},w_{2}\}$ (the gray area). ∎
Figure 2: The $2$-dimensional polytope $\Pi(\mu,\nu)\cap E$ from the proof of 2.28.
2.6 Identifiability from Full Information
Finally, we consider the simplest case where we observe the full OT information (OBS), i.e., the total costs i, OT plans ii and optimal potentials iii.
Theorem 2.29.
Suppose that we observe $K$ pairs of marginals and the corresponding total costs i, OT plans ii and optimal potentials iii. Then, $c$ is identifiable in $\mathbb{R}^{N\times M}$ if and only if for all $(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket$ there exists a $k\in\llbracket K\rrbracket$ such that $\pi^{(k)}_{i,j}>0$. In other words, $c$ is uniquely determined on $\bigcup_{k=1}^{K}\operatorname{supp}\pi^{(k)}$.
Proof.
2.3 implies that the cost matrix $c$ is uniquely determined on $\bigcup_{k=1}^{K}\operatorname{supp}\pi^{(k)}$. Suppose that this support is not full, i.e., there exists a pair $(r,s)\in\llbracket N\rrbracket\times\llbracket M\rrbracket$ with $(r,s)\notin\bigcup_{k=1}^{K}\operatorname{supp}\pi^{(k)}$. Consider the matrix $\tilde{c}\in\mathbb{R}^{N\times M}$ that is equal to $c$ except $\tilde{c}_{r,s}=c_{r,s}+5$. Then, note that $c\leq\tilde{c}$ and thus $(f^{(k)},g^{(k)})\in\Phi_{\tilde{c}}$ for all $k\in\llbracket K\rrbracket$. Hence, 2.3 yields that $\pi^{(k)}$ and $(f^{(k)},g^{(k)})$ are also optimal w.r.t. $\tilde{c}$ for all $k\in\llbracket K\rrbracket$. Consequently, in this case $c$ is not identifiable. ∎
Remark 2.30 (Computation of $c$).
In the setting of Theorem 2.29, the cost matrix $c$ can be directly computed on $\bigcup_{k=1}^{K}\operatorname{supp}\pi^{(k)}$ from the optimal potentials. More specifically, it holds that $c=\max\{f^{(k)}\oplus g^{(k)}:k\in\llbracket K\rrbracket\}$ where the maximum is taken component-wise.
2.7 Identifiability Under Structural Assumptions on the Cost Matrix
We will investigate the impact of structural assumptions on $c$ to identifiability. Suppose that $N=M$. Often, a reasonable assumption is $c_{i,i}=0$, for all $i\in\llbracket N\rrbracket$, which means that the cost of transporting a point to itself is zero. Additionally, we assume that $c$ is symmetric (e.g. when resulting from a metric) and thus belongs to the set
$\operatorname{Sym}_{N}^{0}\coloneqq\{c\in\mathbb{R}^{N\times N}:\ \text{$c_{i,i}=0$ and $c_{i,j}=c_{j,i}$ for all $i,j\in\llbracket N\rrbracket$}\}\,.$ (17)
Denoting with $\bm{{\triangledown}c}$ and $\bm{{\vartriangle}c}$ the vectorized strict upper and lower triangular matrices of $c$, respectively, we have that $c\in\operatorname{Sym}_{N}^{0}$ is fully determined by $\bm{{\triangledown}c}$ (or $\bm{{\vartriangle}c}$). In particular, we only have $(N-1)N/2$ free variables instead of the full $N^{2}$. As such, we need less information to have identifiability of $c$. Clearly our results that rely on the unique solvability of LP(s) (e.g. Theorem 2.6, Theorem 2.13, Theorem 2.21) remain true for this reduced problem. In particular, we obtain the following corresponding result to Theorem 2.22.
Theorem 2.31.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding total costs i and OT plans ii. Let $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$, be as in (RI). Then, the following holds:
  1. a)
    If the system $\{[\bm{{\triangledown}}+\bm{{\vartriangle}}]\bm{u}^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\}$ contains $(N-1)N/2$ linearly independent vectors, then the cost matrix $c$ is identifiable in $\operatorname{Sym}_{N}^{0}$.
  2. b)
    If $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$ for all $k\in\llbracket K\rrbracket$, then the converse holds.
Note that if we only observe the OT plans, we do not have the problem of shift invariance anymore. Indeed, for $a,\,b\in\mathbb{R}^{N}$ it follows that $a\oplus b\in\operatorname{Sym}_{N}^{0}$ if and only if $a\oplus b=0$. Hence, it holds that $\operatorname{Sym}^{0}_{N}=\operatorname{Sym}^{0}_{N}/{\sim_{\oplus}}$. Thus, we obtain the following specialization of Theorem 2.17.
Theorem 2.32.
Suppose that we observe $K$ distinct pairs of marginals and the corresponding OT plans ii. Let $\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}$, $k\in\llbracket K\rrbracket$, be as in (RI). Then, the following holds:
  1. a)
    If it holds for $S\in\mathbb{N}_{0}$ that
    $\dim(\operatorname{span}\{[\bm{{\triangledown}}+\bm{{\vartriangle}}](\bm{\pi}^{(k)}-\bm{u}^{(\ell,k)}):\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\})=(N-1)N/2-S,$
    then, the cost matrix $c$ is identifiable in $\operatorname{Sym}_{N}^{0}$ up to the span of $S$ linearly independent vectors.
  2. b)
    If $\operatorname{coh}\{u^{(\ell,k)}:\ell\in\llbracket L_{k}\rrbracket\}=\operatorname{OTP}_{c}(\pi^{(k)})$ for all $k\in\llbracket K\rrbracket$, then the converse holds.
Proof.
Since $c\in\operatorname{Sym}_{N}^{0}$, it holds for all $x\in\mathbb{R}^{N\times N}$ that
$\langle{c},{x}\rangle=\langle{\bm{{\triangledown}c}},{[\bm{{\triangledown}}+\bm{{\vartriangle}}]\bm{x}}\rangle\,.$
Hence, as in the proof of Theorem 2.17, we conclude that
$\bm{{\triangledown}c}\in\left[\operatorname{span}\left\{[\bm{{\triangledown}}+\bm{{\vartriangle}}](\bm{\pi}^{(k)}-\bm{u}^{(\ell,k)}):\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}\right]^{\perp}\,.$
The rest follows analogously to Theorem 2.17 by considering $c_{\tau}=c+\tau c^{\prime}\in\operatorname{Sym}_{N}^{0}$ where $c^{\prime}\in\operatorname{Sym}_{N}^{0}$ is linearly independent of $c$ and
$\bm{{\triangledown}c^{\prime}}\in\left(\operatorname{span}\left\{[\bm{{\triangledown}}+\bm{{\vartriangle}}](\bm{\pi}^{(k)}-\bm{u}^{(\ell,k)}):\ell\in\llbracket L_{k}\rrbracket,\,k\in\llbracket K\rrbracket\right\}\right)^{\perp}\,.\qed$
3 Statistical Application: Estimation of the Cost
If we receive $NM$ linearly independent OT plans ii and the corresponding total costs i, we would like to solve (15). This can be achieved by computational linear algebra. However, in practice, the values of the OT plans and the total costs may be subjected to some random noise, so that the observation is no longer (15), but
$y^{(k)}\coloneqq\alpha^{(k)}+\varepsilon^{(k)}=\langle{c},{\pi^{(k)}}\rangle+\varepsilon^{(k)}\,,\qquad k\in\llbracket K\rrbracket,$ (18)
where $\{\varepsilon^{(k)}\}_{k=1}^{K}$ is a sequence of independent centered random variables with common variance $\sigma^{2}>0$ (see 3.2 for generalizations). Upon defining the following matrix and vectors
$P_{K}\coloneqq\begin{bmatrix}\bm{\pi}^{(1)}&\ldots&\bm{\pi}^{(K)}\end{bmatrix}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}},\quad\bm{Y}_{K}\coloneqq[y^{(k)}]_{k=1}^{K},\quad\bm{\alpha}_{K}\coloneqq[\alpha^{(k)}]_{k=1}^{K},\quad\bm{\varepsilon}_{K}\coloneqq[\varepsilon^{(k)}]_{k=1}^{K}\,,$
the display (18) can be rewritten as
$P_{K}\bm{c}=\bm{Y}_{K}=\bm{\alpha}_{K}+\bm{\varepsilon}_{K},$ (19)
where $\bm{\varepsilon}_{K}$ is a centered random vector with covariance matrix $\sigma^{2}\operatorname{Id}$. This reveals (19) as a statistical linear model, and we can make use of its well-developed theory. One approach to solve (19) is to employ the least squares estimator (which is statistically efficient when the error is normally distributed):
$\hat{c}_{K}\in\operatorname*{arg\,min}_{c\in\mathbb{R}^{N\times M}}\lVert\bm{Y}_{K}-P_{K}\bm{c}\rVert^{2}_{2}\,.$ (20)
If $NM$ elements of $\{\bm{\pi}^{(k)}\}_{k=1}^{K}$ are linearly independent, the matrix $P_{K}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}P_{K}$ is invertible and the solution can be expressed as
$\hat{\bm{c}}_{K}=(P_{K}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}P_{K})^{-1}P_{K}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\bm{Y}_{K}\,.$
Theorem 3.1.
Assume that there exists a solution $c$ of (15) and that we observe for $K=1,2,\ldots$ a sequence of noisy observations $\bm{Y}_{K}$ in (18). Suppose that there exists a positive definite matrix $Q\in\mathbb{R}^{NM\times NM}$ such that
$\frac{1}{K}P_{K}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}P_{K}\to Q\qquad\text{for }K\to\infty\,.$
Then, the least squares estimator $\hat{c}_{K}$ from (20) is strongly consistent, i.e.,
[…]
If additionally there is a $\delta>0$ such that the $(2+\delta)$-th moment of $\varepsilon^{(k)}$ exists and is uniformly bounded for all $k=1,2,\ldots$, then it follows that
[…]
Proof.
Follows from [White2001, Theorem 2.12 and Exercise 5.12]. ∎
In a similar fashion this can be extended to different estimators such as the $L^{1}$ minimizer of (19), see e.g. [Carroll1982], and different measurement errors $\bm{\varepsilon}_{K}$, e.g., for certain dependency structures, see e.g. [Wu2007].
Remark 3.2.
Our linear model is based on (15) and not (16). If we want to use the additional information provided by (RI), we have a more complicated linear model. Indeed, then the noise is not i.i.d. anymore but coupled, i.e., multiple observations have the exact same noise. This can be dealt with similarly, but now the least squares estimator in (20) has to be replaced by the generalized least squares estimator [Amemiya1985, Rao2008].
3.1 Identifiability in High-Dimensional Linear Models: Sparsity
We again consider the noisy model (18) and assume that $c$ is sparse. In the usual sense, this takes the form of $c_{i,j}=0$ for many pairs of indices $(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket$. As this means that transport between $x_{i}$ and $y_{j}$ is for free and hence often not a reasonable assumption, we utilize a shifted version tailored to OT. More specially, for a predefined value $b_{0}\in\mathbb{R}$, we assume that the set $S\coloneqq\{(i,j)\in\llbracket N\rrbracket\times\llbracket M\rrbracket:c_{i,j}\neq b_{0}\}$ has small cardinality $s\ll NM$, i.e., that the transport cost between many points is equal. In particular, this means that the exact choice of transport between these pairs of points is irrelevant. In this setting, we can make use of this type of sparsity to effectively estimated $c$ even when $NM$ is large compared to the number of observations $K$. We illustrate estimation of $c$ with a corresponding modification of the LASSO estimator [Tibshirani1996]
$\hat{c}_{K}\in\operatorname*{arg\,min}_{c\in\mathbb{R}^{N\times M}}\lVert\bm{Y}_{K}-P_{K}\bm{c}\rVert^{2}_{2}+\lambda_{K}\lVert\bm{c}-b_{0}\rVert_{1}\,.$ (21)
To be able to give error bounds for the LASSO estimator, we require that $P_{K}$ satisfies the restricted eigenvalue condition over $S$ with parameters $\kappa>0$ (cf. [Wainwright2019, Buehlmann2011]). For $x\in\mathbb{R}^{NM}$ denote with $x_{S}$ and $x_{S^{c}}$ the restriction of $x$ to the index sets $S$ or $S^{c}$. Then, $P_{k}$ satisfies the restricted eigenvalue property (REP) over $S$ with parameter $\kappa>0$ if for all $x\in\mathbb{R}^{NM}$ with $\lVert x_{S^{c}}\rVert_{1}\leq 3\lVert x_{S}\rVert_{1}$ it holds that
$\frac{1}{K}\lVert P_{K}x\rVert_{2}^{2}\geq\kappa\lVert x\rVert_{2}^{2}\,.$ (REP)
Naturally, (REP) is satisfied if the observed OT plans contain $\{\delta_{i}\otimes\delta_{j}\mid(i,j)\in S\}$, i.e., $P_{K}$ is the identity matrix when restricting to $S$. This means that we observe all the non-sparse cost coefficients directly (up to noise). However, more generally it is not clear to us what conditions on the transport plans imply (REP).
Under (REP) (and related conditions), error bounds for the LASSO and related estimators have been the topic of a rich body of literature, see e.g. [Giraud2021, Wainwright2019, Buehlmann2011] for recent monographs and further references. Exemplarily, we assume that the noise is normally distributed, then we obtain the following error bound from [Wainwright2019, Example 7.14].
Theorem 3.3.
Assume that $P_{K}$ satisfies (REP) over $S$ with parameter $\kappa>0$ and that the noise $\varepsilon^{(1)},\ldots,\varepsilon^{(K)}$ is i.i.d. zero-mean normally distributed with variance $\sigma^{2}>0$. Then, it holds for any $\delta>0$ for the LASSO type estimator $\hat{\bm{c}}_{K}$ from (21) with regularization parameter
$\lambda_{K}=2\sigma\left[\sqrt{\frac{2\log N+2\log M}{K}}+\delta\right]$
with probability at least $1-2e^{-n\delta^{2}/2}$ that
$\lVert\hat{\bm{c}}_{K}-\bm{c}\rVert_{2}\leq\frac{6\sigma}{\kappa}\sqrt{s}\left[\frac{2\log N+2\log M}{K}+\delta\right]\,.$
Acknowledgements
The research of M. Groppe was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) RTG 2088 “Discovering structure in complex data: Statistics meets Optimization and Inverse Problems” and A. Munk acknowledges support of the DFG CRC 1456 “Mathematics of Experiment” projects A4, C6.
[…]
[title = References]